OpenStudy (anonymous):

Differentiate the function. f(x)= ln[(2x+1)^3/(3x-1)^4]

OpenStudy (mathteacher1729):

Oh wow, you have everything here. Quotient & Chain rule! (or product rule if you want to use that).

OpenStudy (anonymous):

Try the chain rule:\[\frac{da}{db}=\frac{da}{dc} \times\frac{dc}{db}\]

OpenStudy (mathteacher1729):

How far did you get with this problem?

OpenStudy (anonymous):

So I was trying to use the chain rule first but now I am thinking that you are supposed to start with the log and then move to the Quotient rule. Is this correct?

OpenStudy (anonymous):

so could you next write it as ln[(3(2x+1)^2)/(4(3x-1)^3)

OpenStudy (mathteacher1729):

Is this the function? \[\huge f(x)= \ln\left(\frac{(2x+1)^3}{(3x-1)^4}\right)\]

OpenStudy (anonymous):

Yes

OpenStudy (mathteacher1729):

Ok, so the key here is to use log rules to simplify the expression to the sum or difference of logs.

OpenStudy (anonymous):

f(x)= ln[(2x+1)^3/(3x-1)^4] you can write it as.. f(x) = ln(2x+1)^3 - ln(3x-1)^4 then apply derivative...

OpenStudy (mathteacher1729):

http://www.purplemath.com/modules/logrules.htm

OpenStudy (anonymous):

@AJW99 got it??

OpenStudy (anonymous):

Still trying to wrap my brain around it

OpenStudy (mathteacher1729):

First log rule you'll use here... \[\huge \ln\left(\frac{a}{b}\right) = \ln(a) - \ln(b)\]

OpenStudy (mathteacher1729):

In this case a = (2x+1)^3 and b = (3x-1)^4

OpenStudy (anonymous):

okay so then it would look like this correct? ln(2x+1)^3 - ln(3x-1)^4

OpenStudy (anonymous):

so next would you use the rule d/dx [f(x)]^b=b[f(x)]^b-1 f'(x)?

OpenStudy (anonymous):

I am stuck.

OpenStudy (anonymous):

Can you still help me I have a really hard time with implicit differentiation

OpenStudy (anonymous):

sure..

OpenStudy (anonymous):

@AJW99 can you do it now... or i'll solve it till the end??

OpenStudy (anonymous):

Thank you so much for all of your help! I really appreciate the help!

OpenStudy (anonymous):

no problem... :)

OpenStudy (mathteacher1729):

There is a simpler way about it... Once you have \[\huge \ln((2x+1)^2)-\ln((3x-1)^4)\] you can use another log rule... \[\huge \ln((a)^b)=b\cdot\ln(a)\] In other words, \[\huge f(x) = 2\cdot \ln(2x+1)-4\cdot \ln(3x-1)\] Differentiating this is much easier, in my opinion. :)

OpenStudy (anonymous):

@mathteacher1729 I think I have done the same way...

OpenStudy (mathteacher1729):

@Kashan Yes, our final answers agree. But you differentiated using the exponent rule and then in your final result did simplification to yield the same answer. This solution (simplifying as much as possible using logs) avoids having to do any canceling at the end. Both are correct, though. :)

OpenStudy (anonymous):

Yes I agree.. thanks :)